# Toward Universal Decoding of Binary Linear Block Codes via Enhanced Polar Transformations

### v1

@@ Abstract
Binary linear block codes (BLBCs) are essential to modern communication, but their diverse structures often require multiple decoders, increasing complexity. This work introduces enhanced polar decoding (${\mathsf{PD}^{+}}$), a universal soft decoding algorithm that transforms any BLBC into a polar-like code compatible with efficient polar code decoders such as successive cancellation list (SCL) decoding. Key innovations in ${\mathsf{PD}^{+}}$ include pruning polar kernels, shortening codes, and leveraging a simulated annealing algorithm to optimize transformations. These enable ${\mathsf{PD}^{+}}$ to achieve competitive or superior performance to state-of-the-art algorithms like OSD and GRAND across various codes, including extended BCH, extended Golay, and binary quadratic residue codes, with significantly lower complexity. Moreover, ${\mathsf{PD}^{+}}$ is designed to be forward-compatible with advancements in polar code decoding techniques and AI-driven search methods, making it a robust and versatile solution for universal BLBC decoding in both present and future systems.
@@ I Introduction
Binary linear block codes (BLBC), a significant subclass of error correction codes, have been widely used in modern communication systems [1]. Since Hamming’s pioneering work [2], numerous types of codes have been developed, each with unique strengths. For instance, BCH codes [1] are renowned for their large minimum Hamming distance, low-density parity-check (LDPC) codes [3] enable low-complexity soft decoding and provide excellent finite-length performance, and polar codes [4] are celebrated for being asymptotically capacity-achieving and free from error floors [5]. The diverse strengths of these codes have led to the implementation of multiple types of codes within a single system. A recent example is the 5G cellular system [6, 7], where polar codes and LDPC codes are employed as the coding techniques for the control and data channels, respectively. To decode those codes, the current approach is to equip each device with multiple decoders, one for each type of code, which necessarily increases the decoding complexity and the decoder size significantly. This motivates the pursuit of universal decoding; that is, a low-complexity decoding algorithm that can decode all types of codes.
Another compelling motivation for pursuing universal decoding is the challenge of decoding powerful algebraic codes. Algebraic codes form a substantial subset of BLBC, utilizing algebraic structures to achieve excellent Hamming distances and robust error-correction capabilities [8]. For instance, BCH codes are widely employed in disk storage, CDs, and satellite communications due to their strong multi-error correction capability. Similarly, Golay codes, small perfect error-correcting codes, are used in applications like deep space and radio communications. However, despite their many advantages, most algebraic codes face a significant limitation: they are inherently difficult to soft-decode [1]. [^1] While soft-decoding methods for algebraic codes do exist, see for example [9, 10], they remain complex and are not easily generalizable to other codes. As a result, efficient and low-complexity soft-decoding algorithms for algebraic codes are still in demand. Developing a low-complexity universal soft-decoding approach could effectively address this challenge.
@@
In [14], we proposed a universal decoding algorithm named polar decoding (${\mathsf{PD}}$). The core idea of ${\mathsf{PD}}$ is to transform a BLBC into a polar code with dynamic frozen bits [15] and decode it using a polar code decoding algorithm. This is of enormous practical importance, as the implementation of decoding algorithms for polar codes has become highly mature [16, 17, 18, 19], thanks to their adoption in 5G. The transformation is achieved through a permutation operation, where different permutations result in polar codes with different sets of dynamic frozen bits. A brute-force search was then used to identify a permutation that yields good decoding performance. In [14], we demonstrated that ${\mathsf{PD}}$ achieves near-MLD performance with lower decoding complexity than OSD for various extended BCH codes and an extended Golay code [20]. However, a challenging case was also identified in [14], pointing out the limitations of ${\mathsf{PD}}$. Specifically, the following generator matrix was considered:
@@
  • In ${\mathsf{PD}}$, the set of kernels (multi-kernel polar codes [22] are allowed) for the polar codes is first determined, and then a suitable permutation matrix is searched for to transform the target BLBC into a polar code with that set of kernels. This restricts the transformed code to the strict family of polar codes whose generator matrices admit a Kronecker product structure. Moreover, it is challenging to decide which set of kernels to start with and in what order to arrange them. For example, for a code of length 24 that can be transformed into polar codes with kernels of sizes $3\times 2\times 2\times 2$, it is difficult to determine where to place the kernel of size 3 and which kernel of size 3 to adopt. In ${\mathsf{PD}^{+}}$, we propose a novel transformation that converts a BLBC into a polar code with a pruned kernel. In this approach, we always begin with a polar code using Arikan’s kernel and remove some edges to form a polar code with pruned kernels. The question of which edges to prune is then formulated as a local search problem.
  • Originally, in ${\mathsf{PD}}$, the multi-kernel technique is adopted to control the code length. However, for ${\mathsf{PD}^{+}}$, as mentioned above, we adopt the pruning technique to modify Arikan’s kernel. Consequently, another technique is required to control the code length. The second ingredient in the proposed ${\mathsf{PD}^{+}}$ is to further shorten the transformed polar code, tailoring it to meet the desired code length. The problem of determining which bits to shorten is again formulated as a local search problem.
@@
In summary, the main contribution of this work is the proposal of a novel universal decoding algorithm, ${\mathsf{PD}^{+}}$, which transforms a BLBC into a polar-like code. After transformation, this code becomes a shortened version of a polar code with pruned Arikan’s kernels and dynamic frozen bits; thereby, can still be decoded by a decoding algorithm for polar codes. The algorithm is universal in the sense that it can decode any BLBC, while its performance depends on the quality of the search results. Extensive simulations presented in this paper demonstrate that, for many codes—including extended BCH codes, extended Golay codes, and binary quadratic residue codes—${\mathsf{PD}^{+}}$ together with SCL decoding achieves performance that is better than or comparable to OSD, GRAND, and ${\mathsf{PD}}$, while offering significantly lower decoding complexity.
@@ II-A Decoding of BLBC
Let $\mathcal{C}$ be an $(n,k)$ BLBC that encodes a $k$-bit message $\mathbf{m}\in\mathbb{F}_{2}^{k}$ into a $n$-bit codeword $\mathbf{c}\in\mathbb{F}_{2}^{n}$ whose relationship is given by
@@
where $\mathbf{G}\in\mathbb{F}_{2}^{k\times n}$ is a generator matrix. A modulated codeword $\mathbf{x}$ is sent to the channel $\mathcal{W}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ and a noisy output $\mathbf{y}$ arrives at the receiver. A decoding function $\mathsf{dec}$ is applied on $\mathbf{r}$ a post-processed version of $\mathbf{y}$ to an estimate $\hat{\mathbf{m}}=\mathsf{dec}(\mathbf{r})$. [^2] Depending on the considered scenario, the modulation and post-processing could be omitted. For example, for binary symmetric channels, the codeword $\mathbf{c}$ is directly sent and the received word $\mathbf{y}$ is directly fed into $\mathsf{dec}$. The probability of (frame) error is $p_{e}=\mathbb{P}(\hat{\mathbf{m}}\neq\mathbf{m})$. The goal of this paper is to design a $\mathsf{dec}$ that can provide low $p_{e}$ for a large class of codes $\mathcal{C}$.
@@ II-B Conventional Polar Codes
Polar codes, introduced by Erdal Arikan in 2009 [4], represent the first family of capacity-achieving codes for binary memoryless channels, featuring explicit construction and low encoding/decoding complexity. What follows is a brief review of polar codes.
@@
We start by explaining the channel polarization for $n=2$. Consider the basic polarization kernel (or Arikan’s kernel) as shown in Fig. 1, which transforms $(u_{1},u_{2})$ into $(x_{1},x_{2})$ by $x_{1}=u_{1}\oplus u_{2}$ and $x_{2}=u_{2}$. i.e., $\mathbf{x}=\mathbf{u}\mathbf{G}_{2}$, where
@@
Figure 1: Arikan’s kernel.
At the receiver, two new channels are formed. For the first one, we decode $u_{1}$ by $(y_{1},y_{2})$, giving rise to ${\mathcal{W}}^{-}:{\mathcal{X}}\rightarrow{\mathcal{Y}}^{2}$, which we abbreviated by the operation $\boxplus$. i.e., ${\mathcal{W}}^{-}={\mathcal{W}}\boxplus{\mathcal{W}}$. For the second one, we decode $u_{2}$ by $(u_{1},y_{1},y_{2})$, giving rise to ${\mathcal{W}}^{+}:{\mathcal{X}}\rightarrow{\mathcal{Y}}^{2}\times{\mathcal{X}}$, represented by the operation $\boxtimes$. Thus, ${\mathcal{W}}^{+}={\mathcal{W}}\boxtimes{\mathcal{W}}$. Their transition probabilities are given as follows:
$({\mathcal{W}}\boxplus{\mathcal{W}})(y_{1},y_{2}|u_{1})=\frac{1}{2}\sum_{u_{2}\in{\mathcal{X}}}{\mathcal{W}}(y_{1}|u_{1}\oplus u_{2}){\mathcal{W}}(y_{2}|u_{2}),$ (5)
@@
$({\mathcal{W}}\boxtimes{\mathcal{W}})(y_{1},y_{2},u_{1}|u_{2})=\frac{1}{2}{\mathcal{W}}(y_{1}|u_{1}\oplus u_{2}){\mathcal{W}}(y_{2}|u_{2}),$ (6)
@@
$\hskip 10.0pt\cdot\sum_{u_{2i+1,o}^{2N}}\frac{1}{2^{N-1}}{\mathcal{W}}^{N}(y_{1}^{N}|u_{1,o}^{2N}\oplus u_{1,e}^{2N}),$ (10)
@@
${\mathcal{W}}^{2N}_{2i}(y_{1}^{2N},u_{1}^{2i-1}|u_{2i})={\mathcal{W}}^{N}\boxtimes{\mathcal{W}}^{N}$
@@
$\hskip 10.0pt\cdot\sum_{u_{2i+1,o}^{2N}}\frac{1}{2^{N-1}}{\mathcal{W}}^{N}(y_{1}^{N}|u_{1,o}^{2N}\oplus u_{1,e}^{2N}).$ (11)
@@
Figure 2: LHS: Recursive structure of Arikan’s polarization with $n=8$. RHS: The bit-reversed version of the same structure.
@@ II-B1 Other Polarization Kernel and Multi-Kernel Polar Codes
The asymptotic performance of polar codes has been shown to be tightly connected to the polarization kernel, which is the object undergoing the Kronecker product. For instance, Arikan’s kernel, as shown in (4), is a size-2 kernel. Kernels with better asymptotic performance than Arikan’s kernel have also been developed; see, for example, [30, 31]. Instead of having a length of $2^{m}$, a polar code with a kernel of size $q$ will have a length of $q^{m}$.
@@ II-C Review of Polar Decoding
A $(n,k)$ BLBC can be transformed into a (possibly multi-kernel) polar code with dynamic frozen bits.
@@
$\mathcal{C}\triangleq$$\left\{\mathbf{c}=\mathbf{m}(\mathbf{E}^{-1}\mathbf{E})\mathbf{G}(\mathbf{P}^{-1}\mathbf{G}_{n}^{-1}\mathbf{G}_{n}\mathbf{P})|\mathbf{m}\in\mathbb{F}_{2}^{k}\right\},$
$=$ $\left\{\mathbf{c}=(\mathbf{m}\mathbf{E}^{-1})(\mathbf{E}\mathbf{G}\mathbf{P}^{-1}\mathbf{G}_{n}^{-1})\mathbf{G}_{n}\mathbf{P}|\mathbf{m}\in\mathbb{F}_{2}^{k}\right\}$
$=$ $\left\{\mathbf{c}=\mathbf{c}_{p}\mathbf{P}|\mathbf{c}_{p}\in\mathcal{C}_{p}\right\},$ (15)
where $\mathbf{G}_{n}$ is the generator matrix of a polar code of size $n$, $\mathbf{P}$ is an $n\times n$ permutation matrix, and $\mathbf{E}$ is the elimination matrix in Gaussian elimination procedure that transforms $\mathbf{G}\mathbf{P}^{-1}\mathbf{G}_{n}^{-1}$ into upper-trapezoidal form. In ${\mathsf{PD}}$, by permuting the received signal $\mathbf{y}$ with $\mathbf{P}^{-1}$, the decoding problem [^4] Since $\mathbf{E}^{-1}$ is always non-singular, decoding $\mathbf{m}$ and $\mathbf{m}\mathbf{E}^{-1}$ are equivalent. reduces to that of a polar code with dynamic frozen bits induced by $\mathbf{M}_{\rm DF}=\mathbf{E}\mathbf{G}\mathbf{P}^{-1}\mathbf{G}_{n}^{-1}$.
@@ III-A1 Challenging case
In [14], a challenging case was identified, as shown in (1). It was demonstrated that when $\mathcal{W}$ is an AWGN channel with a raw bit error rate of $0.01$, the sum of the Bhattacharyya parameters in (13) amounts to $0.198997\times 3=0.596991$. However, among the $8!=40320$ possible choices of permutation matrices, the minimum achievable sum of Bhattacharyya parameters is $0.830539+0.149237+0.003134=0.98291$, which is significantly larger than $0.596991$. This result highlights that ${\mathsf{PD}}$ is fundamentally inadequate in addressing such cases, and the issue cannot simply be attributed to the large search space.

### v2

@@ Abstract
Binary linear block codes (BLBCs) are essential to modern communication, but their diverse structures often require tailor-made decoders, increasing complexity. This work introduces enhanced polar decoding (${\mathsf{PD}^{+}}$), a universal soft decoding algorithm that transforms any BLBC into a polar-like code compatible with efficient polar code decoders such as successive cancellation list (SCL) decoding. Key innovations in ${\mathsf{PD}^{+}}$ include pruning polar kernels, shortening codes, and leveraging a simulated annealing algorithm to optimize transformations. These enable ${\mathsf{PD}^{+}}$ to achieve competitive or superior performance to state-of-the-art algorithms like OSD and GRAND across various codes, including extended BCH, extended Golay, and binary quadratic residue codes, with significantly lower complexity. Moreover, ${\mathsf{PD}^{+}}$ is designed to be forward-compatible with advancements in polar code decoding techniques and AI-driven search methods, making it a robust and versatile solution for universal BLBC decoding in both present and future systems.
@@ I Introduction
Binary linear block codes (BLBC), a significant subclass of error correction codes, have been widely used in modern communication systems [1]. Since Hamming’s pioneering work [2], numerous types of codes have been developed, each with unique strengths. For instance, BCH codes [1] are renowned for their large minimum Hamming distance, low-density parity-check (LDPC) codes [3] enable low-complexity soft decoding and provide excellent finite-length performance, and polar codes [4] are celebrated for being asymptotically capacity-achieving and free from error floors [5]. The diverse strengths of these codes have led to the implementation of multiple types of codes within a single system. A recent example is the 5G cellular system [6, 7], where polar codes and LDPC codes are employed as the coding techniques for the control and data channels, respectively. To decode those codes, the current approach is to equip each device with multiple decoders, one for each type of code, which necessarily increases the decoding complexity and decoder size. This motivates the pursuit of universal decoding, namely, a low-complexity decoding algorithm capable of decoding all types of codes. In particular, we consider a parametrized universal decoder, which enables multiple codes to share the same decoding circuit; only a small set of parameters needs to be stored and applied, greatly reducing hardware overhead.
Another compelling motivation for pursuing universal decoding is the challenge of decoding powerful algebraic codes. Algebraic codes form a substantial subset of BLBC, utilizing algebraic structures to achieve excellent Hamming distances and robust error-correction capabilities [8]. For instance, BCH codes are widely employed in disk storage, CDs, and satellite communications due to their strong multi-error correction capability. Similarly, Golay codes, small perfect error-correcting codes, are used in applications like deep space and radio communications. However, despite their many advantages, most algebraic codes face a limitation: they are inherently difficult to soft-decode [1]. [^1] While soft-decoding methods for algebraic codes do exist, see for example [9, 10], they remain complex and are not easily generalizable to other codes. As a result, efficient and low-complexity soft-decoding algorithms for algebraic codes are still in demand. Developing a low-complexity universal soft-decoding approach could effectively address this challenge.
@@
In [14], we proposed a universal decoding algorithm named polar decoding (${\mathsf{PD}}$). The core idea of ${\mathsf{PD}}$ is to transform a BLBC into a polar code with dynamic frozen bits [15] and decode it using a polar code decoding algorithm. This is of enormous practical importance, as the implementation of decoding algorithms for polar codes has become highly mature [16, 17, 18, 19], thanks to their adoption in 5G. The transformation is achieved through a permutation operation, where different permutations result in polar codes with different sets of dynamic frozen bits. A brute-force search was then used to identify a permutation that yields good decoding performance. In [14], we demonstrated that ${\mathsf{PD}}$ achieves near-MLD performance with lower decoding complexity than OSD for various extended BCH (eBCH) codes and an extended Golay (eGolay) code [20]. However, a challenging case was also identified in [14], pointing out the limitations of ${\mathsf{PD}}$. Specifically, the following generator matrix was considered:
@@
  • In ${\mathsf{PD}}$, the set of kernels (multi-kernel polar codes [22] are allowed) for the polar codes is first determined, and then a suitable permutation matrix is searched for to transform the target BLBC into a polar code with that set of kernels. This restricts the transformed code to the strict family of polar codes whose generator matrices admit a Kronecker product structure. Moreover, it is challenging to decide which set of kernels to start with and in what order to arrange them. For example, for a code of length 24 that can be transformed into polar codes with kernels of sizes $3\times 2\times 2\times 2$, it is difficult to determine where to place the kernel of size 3 and which kernel of size 3 to adopt. In ${\mathsf{PD}^{+}}$, we propose a novel transformation that converts a BLBC into a polar code with a pruned kernel. In this approach, we always begin with a polar code using Arıkan’s kernel and remove some edges to form a polar code with pruned kernels. The question of which edges to prune is then formulated as a local search problem.
  • Originally, in ${\mathsf{PD}}$, the multi-kernel technique is adopted to control the code length. However, for ${\mathsf{PD}^{+}}$, as mentioned above, we adopt the pruning technique to modify Arıkan’s kernel. Consequently, another technique is required to control the code length. The second ingredient in the proposed ${\mathsf{PD}^{+}}$ is to further shorten the transformed polar code, tailoring it to meet the desired code length. The problem of determining which bits to shorten is again formulated as a local search problem.
@@
In summary, the main contribution of this work is the proposal of a novel universal decoding algorithm, ${\mathsf{PD}^{+}}$, which transforms a BLBC into a polar-like code. After transformation, this code becomes a shortened version of a polar code with pruned Arıkan’s kernels and dynamic frozen bits. As a result, it can still be decoded by a decoding algorithm for polar codes. The algorithm is universal in the sense that it can decode any BLBC, while its performance depends on the quality of the search results. Extensive simulations presented in this paper demonstrate that, for many codes—including eBCH codes, eGolay codes, and binary quadratic residue codes—${\mathsf{PD}^{+}}$ together with SCL decoding achieves performance that is better than or comparable to OSD, GRAND, and ${\mathsf{PD}}$, while offering significantly lower decoding complexity.
@@ II-A Decoding of BLBC
Let $\mathcal{C}$ be an $(n,k)$-BLBC that encodes a $k$-bit message $\mathbf{m}\in\mathbb{F}_{2}^{k}$ into a $n$-bit codeword $\mathbf{c}\in\mathbb{F}_{2}^{n}$ whose relationship is given by
@@
where $\mathbf{G}\in\mathbb{F}_{2}^{k\times n}$ is a generator matrix. A modulated codeword $\mathbf{x}$ is sent to the channel $\mathcal{W}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ and a noisy output $\mathbf{y}$ arrives at the receiver. A decoding function $\mathsf{dec}$ is applied on $\mathbf{r}$ a post-processed version of $\mathbf{y}$ to an estimate $\hat{\mathbf{m}}=\mathsf{dec}(\mathbf{r})$. The goal of this paper is to design a $\mathsf{dec}$ that can provide low $p_{e}$ for a large class of codes $\mathcal{C}$.
@@ II-B Conventional Polar Codes
Polar codes, introduced by Erdal Arıkan in 2009 [4], represent the first family of capacity-achieving codes for binary memoryless channels, featuring explicit construction and low encoding/decoding complexity. What follows is a brief review of polar codes.
@@
We start by explaining the channel polarization for $n=2$. Consider the basic polarization kernel (or Arıkan’s kernel) as shown in Fig. 1, which transforms $(u_{1},u_{2})$ into $(x_{1},x_{2})$ by $x_{1}=u_{1}\oplus u_{2}$ and $x_{2}=u_{2}$. i.e., $\mathbf{x}=\mathbf{u}\mathbf{G}_{2}$, where
@@
Figure 1: Arıkan’s kernel.
[…]
@@
$({\mathcal{W}}\circledast{\mathcal{W}})(y_{1},y_{2},u_{1}|u_{2})=\frac{1}{2}{\mathcal{W}}(y_{1}|u_{1}\oplus u_{2}){\mathcal{W}}(y_{2}|u_{2}),$ (6)
@@
×$\hskip 10.0pt\times\sum_{u_{2i+1,o}^{2N}}\frac{1}{2^{N-1}}{\mathcal{W}}^{N}(y_{1}^{N}|u_{1,o}^{2N}\oplus u_{1,e}^{2N}),$ (10)
@@
${\mathcal{W}}^{2N}_{2i}(y_{1}^{2N},u_{1}^{2i-1}|u_{2i})={\mathcal{W}}^{N}\circledast{\mathcal{W}}^{N}$
@@
×$\hskip 10.0pt\times\sum_{u_{2i+1,o}^{2N}}\frac{1}{2^{N-1}}{\mathcal{W}}^{N}(y_{1}^{N}|u_{1,o}^{2N}\oplus u_{1,e}^{2N}).$ (11)
@@
Figure 2: LHS: Recursive structure of Arıkan’s polarization with $n=8$. RHS: The bit-reversed version of the same structure.
@@ II-B1 Other Polarization Kernel and Multi-Kernel Polar Codes
The asymptotic performance of polar codes has been shown to be tightly connected to the polarization kernel, which is the object undergoing the Kronecker product. For instance, Arıkan’s kernel, as shown in (4), is a size-2 kernel. Kernels with better asymptotic performance than Arıkan’s kernel have also been developed; see, for example, [30, 31]. Instead of having a length of $2^{m}$, a polar code with a kernel of size $q$ will have a length of $q^{m}$.
@@ II-C Review of Polar Decoding
An $(n,k)$-BLBC can be transformed into a (possibly multi-kernel) polar code with dynamic frozen bits.
@@
$\mathcal{C}\triangleq$$\left\{\mathbf{c}=\mathbf{m}(\mathbf{E}^{-1}\mathbf{E})\mathbf{G}(\mathbf{P}^{-1}\mathbf{G}_{n}^{-1}\mathbf{G}_{n}\mathbf{P})\mid\mathbf{m}\in\mathbb{F}_{2}^{k}\right\},$
$=$ $\left\{\mathbf{c}=(\mathbf{m}\mathbf{E}^{-1})(\mathbf{E}\mathbf{G}\mathbf{P}^{-1}\mathbf{G}_{n}^{-1})\mathbf{G}_{n}\mathbf{P}\mid\mathbf{m}\in\mathbb{F}_{2}^{k}\right\}$
$=$ $\left\{\mathbf{c}=\mathbf{c}_{p}\mathbf{P}\mid\mathbf{c}_{p}\in\mathcal{C}_{p}\right\},$ (15)
where $\mathbf{G}_{n}$ is the generator matrix of a polar code of size $n$, $\mathbf{P}$ is an $n\times n$ permutation matrix, $\mathcal{C}_{p}$ is the polar code corresponding to $\mathbf{G}_{n}$ and dynmamic fronzen bits induced by $\mathbf{M}_{\rm DF}=\mathbf{E}\mathbf{G}\mathbf{P}^{-1}\mathbf{G}_{n}^{-1}$, and $\mathbf{E}$ is the elimination matrix in Gaussian elimination procedure that transforms $\mathbf{G}\mathbf{P}^{-1}\mathbf{G}_{n}^{-1}$ into upper-trapezoidal form. In ${\mathsf{PD}}$, by permuting the received signal $\mathbf{y}$ with $\mathbf{P}^{-1}$, the decoding problem [^3] Since $\mathbf{E}^{-1}$ is always non-singular, decoding $\mathbf{m}$ and $\mathbf{m}\mathbf{E}^{-1}$ are equivalent. reduces to that of a polar code with dynamic frozen bits induced by $\mathbf{M}_{\rm DF}$.
@@ III-A1 Challenging case
In [14], a challenging case was identified, as shown in (1). It was demonstrated that when $\mathcal{W}$ is an AWGN channel with a raw bit error rate of $0.01$, the sum of the Bhattacharyya parameters in (13) amounts to $0.198997\times 3=0.596991$. However, among the $8!=40320$ possible choices of permutation matrices, the minimum achievable sum of Bhattacharyya parameters is $0.830539+0.149237+0.000002=0.979778$, which is significantly larger than $0.596991$. This result highlights that ${\mathsf{PD}}$ is fundamentally inadequate in addressing such cases, and the issue cannot simply be attributed to the large search space.